# UnMarker: A Universal Attack on Defensive Watermarking

### v1

@@ 5.1. Attacking Non-Semantic Watermarks
Final High-Frequency Destruction Procedure. (for non-semantic watermarks) Using the above components, we implement eq. (1). We only need to identify the rules for updating the sample to define our search space. Since we seek high-frequency disruptions that affect the rates at which individual pixels vary w.r.t. their neighbors (as opposed to collective low-frequency changes), we should modify each pixel directly at each step. Thus, our algorithm accepts watermarked image $x_{w}$, LPIPS loss $\ell_{p}$, LPIPS threshold $t_{\ell_{p}}$, and $\|\|_{2}$ threshold $t_{\|\|_{2}}$, and finds modifier $\delta$ that maximizes DFL as follows:
@@
5.2. Attacking Semantic Watermarks
Semantic watermarks reside in low frequencies, making them difficult to address. In addition to their large magnitudes (see §4.1.2), low frequencies correspond to many neighboring pixels varying gradually, indicating these watermarks are rooted in the image’s objects’ consistencies. Thus, removing them requires altering pixels systemically via structural changes to the different regions. Direct optimizations similar to §5.1 fail since they operate on the image and, at each step, individual pixels are updated in different directions. This targets higher frequencies, introducing significant changes yet only to a few pixels, failing to progress further due to the perceptual similarity constraint and incurring minimal changes at low frequencies. We could limit the procedure to optimize DFL over low-frequencies only. Yet, as significant changes are required in this range, without a methodical scheme, we will converge to a sub-optimal solution wherein a certain region receives substantial modifications, leading the perceptual loss to prevent further updates. This is because the process fails to account for inter-pixel links, and DFL does not spread the changes throughout the image.
(a) Basic: To calculate the output at pixel $(i,\;j)$, the patch $PE^{i,\;j}_{\tiny\mbox{$M{\times}N$}}$, where $M{\times}N$ is the kernel size, around $(i,\;j)$ is element-wise multiplied by the spectral kernel $\mathcal{K}$, summing over this product.
(b) Guided: The color kernel $\mathcal{C}^{\mathcal{I},\sigma_{c}}_{i,\;j}$ is first computed, assigning a modulation factor to each neighbor in the vicinity $PE^{i,\;j}_{\tiny\mbox{$M{\times}N$}}$ of $(i,\;j)$. These are element-wise multiplied by the spectral kernel $\mathcal{K}$ to limit the contribution from neighbors whose values differ significantly from $(i,\;j)$’s. This yields the effective kernel $\mathcal{V}^{\mathcal{I},\mathcal{K},\sigma_{c}}_{i,\;j}$ used to calculate the output at $(i,\;j)$ similar to basic filters.
Figure 3. Operation of known filters.
We develop a visually-aware attack inspired by image filtering. While filters apply systemic spectral disruptions, they fail to remove watermarks due to their content-agnostic nature, limiting the distortions they may introduce without leaving crucial visual cues. Our adversarial alternative retains quality by learning the filters’ weights to solve an optimization problem that maximizes low-frequency disruptions, resorting to a perceptual loss to preserve critical information. First, we motivate our design through guiding principles. We then provide technical details and present our optimizable filters. Finally, we introduce our system’s components.
5.2.1. Motivation
Convolution in the spatial domain translates to multiplication in the frequency space: Given the spectral representations of the image and filter, it produces an image with a spectrum equivalent to their product. Thus, filters are good candidates for spectral manipulations. Yet, the outputs at any pixel are calculated using the same kernel. Hence, their potential to remove watermarks is limited (Lukas and Kerschbaum, 2023; Zhao et al., 2023c) due to their content-agnostic nature, making them only applicable with mild parameters as stronger blurring impacts quality.
Guided Filtering: These advanced filters tackle the information loss under stronger distortions, allowing more effective noise reduction while preserving crucial data. They are edge-preserving, minimizing the distortion at these critical boundary points between different objects. Instead of using a uniform kernel, the process involves an additional per-pixel color kernel that weakens the filter at edge points based on geometric principles: If some values in the pixel’s vicinity drastically differ from it, then it lies on an edge, as these neighbors belong to other components and the spectral kernel’s strength at this point must be diminished to preserve defining information. The color kernel computes a modulation factor for every neighbor in the filter’s vicinity around $(i,\;j)$ based on its value’s difference from $(i,\;j)$’s. The effective kernel for calculating $(i,\;j)$’s output is the Hadamard product of the two kernels— see Fig. 3(b). Guided filters fail to remove watermarks despite allowing for stronger blurring (see §6.2.1) since they merely follow geometric heuristics that fail to fully capture human vision or target specific image characteristics, limiting the allowed distortion. In practice, less visually-critical areas can be changed drastically, regardless.
5.2.2. Formalizing Filters
Given image $\mathcal{I}\in\mathbb{R}^{{\tiny{\mbox{$Q\times R$}}}}$, $PE^{i,\;j}_{\tiny\mbox{$M{\times}N$}}:\mathbb{R}^{{\tiny{\mbox{$Q\times R$}}}}$ $\longrightarrow\mathbb{R}^{{\tiny{\mbox{$M\times N$}}}}$ denotes the local patch extractor that outputs the $M{\times}N$ vicinity around its $(i,\;j)^{th}$ pixel s.t. this pixel is at the center of this patch (i.e., $PE^{i,\;j}_{\tiny\mbox{$M{\times}N$}}(\mathcal{I})({\small\mbox{$\left\lfloor\frac{M}{2}\right\rfloor$}},\;{\small\mbox{$\left\lfloor\frac{N}{2}\right\rfloor$}})=\mathcal{I}(i,\;j)$)— see Fig. 3(a). The global patch extractor $PE_{\tiny\mbox{$M{\times}N$}}:\mathbb{R}^{{\tiny{\mbox{$Q\times R$}}}}\longrightarrow\mathbb{R}^{{\tiny{\mbox{$Q\times R\times M\times N$}}}}$ outputs a matrix storing $PE^{i,\;j}_{\tiny\mbox{$M{\times}N$}}$ for every $(i,\;j)$. That is, $PE_{\tiny\mbox{$M{\times}N$}}(\mathcal{I})$’s $(i,\;j)^{th}$ cell contains the $M{\times}N$ patch around the $(i,\;j)^{th}$ pixel. We now formalize different filters to present our solution.
Basic Filtering: Basic filters simply apply the known convolution with a fixed spectral kernel $\mathcal{K}\in\mathbb{R}^{{\tiny{\mbox{$M\times N$}}}}$. Denoting element-wise multiplication as $\odot$, these filters can be written as $\mathcal{F}_{\mathcal{K}}:\mathbb{R}^{{\tiny{\mbox{$Q\times R$}}}}\longrightarrow\mathbb{R}^{{\tiny{\mbox{$Q\times R$}}}}$ s.t. $\mathcal{F}_{\mathcal{K}}(I)(i,\;j)=\sum_{m,n}PE^{i,\;j}_{\tiny\mbox{$M{\times}N$}}(I)\odot\mathcal{K}$ (see Fig. 3(a)).
Guided Filtering: These employ additional per-pixel color kernels to modulate distortions at critical points (Fig. 3(b)). Given kernel size $M{\times}N$, the color kernel for each $(i,\;j)$ is the matrix whose $(m,\;n)$ cell is $\mathcal{C}^{\mathcal{I},\sigma_{c}}_{i,\;j}(m,\;n)$ = $e^{\mbox{\small${-\frac{\|\mathcal{I}(i+m-\left\lfloor\frac{M}{2}\right\rfloor,\;j+n-\left\lfloor\frac{N}{2}\right\rfloor)-\mathcal{I}(i,\;j)\|_{1}}{2\sigma_{\mathcal{C}}^{2}}}$}}$. $\sigma_{c}$ is a permissiveness controlling the modulation intensity. The $\|\|_{1}$ norm accounts for each pixel potentially having multiple color channels. For instance, RGB images associate each pixel with three values. The effective kernel for computing the output at $(i\;j)$ is then $\mathcal{V}^{\mathcal{I},\mathcal{K},\sigma_{c}}_{i,\;j}={\small\mbox{$\frac{\mathcal{C}^{\mathcal{I},\sigma_{c}}_{i,\;j}\odot\mathcal{K}}{\underset{m,n}{\large\mbox{$\sum$}}\mathcal{C}^{\mathcal{I},\sigma_{c}}_{i,\;j}\odot\mathcal{K}}$}}$. At the center, $\mathcal{C}^{\mathcal{I},\sigma_{c}}_{i,\;j}(\left\lfloor\frac{M}{2}\right\rfloor,\;\left\lfloor\frac{N}{2}\right\rfloor)=1$, indicating the contribution of the original $\mathcal{I}(i,\;j)$ is not decayed. For neighbors, large distances decrease their coefficients and limit their contributions. By defining $\mathcal{V}^{I,\mathcal{K},\sigma_{c}}$ as the matrix whose $(i,\;j)^{th}$ cell contains $\mathcal{V}^{\mathcal{I},\mathcal{K},\sigma_{c}}_{i,j}$, the guided filter is $GF_{\mathcal{K},\sigma_{c}}(\mathcal{I})=\underset{m,n}{\large\mbox{$\sum$}}PE_{\tiny\mbox{$M{\times}N$}}(\mathcal{I})\odot\mathcal{V}^{\mathcal{I},\mathcal{K},\sigma_{c}}$.
Optimizable Filters (Ours): The drawback of guided filtering is that it employs the same spectral kernel $\mathcal{K}$ for all pixels, changing only the color kernel $\mathcal{C}^{\mathcal{I},\sigma_{c}}_{i,\;j}$, computed via the same formula at all locations. Instead, we allow each pixel to have a unique spectral kernel ${\mathcal{K}_{{}_{{}_{\;\;i,j}}}}$, with weights that can differ significantly between locations based on feedback from a perceptual metric that determines their visual importance and allows maximizing the destruction at less critical regions. Thus, our filters consist of a set $\mathcal{K}^{*}$ of local ${\mathcal{K}_{{}_{{}_{\;\;i,j}}}}$’s for each $(i,\;j)$. We include color kernels to ensure the optimization does not explore unwanted directions that violate vital geometric constraints. While these may be limiting, ignoring them causes unnatural distortions if the differences between neighboring pixels are extremely large, making them essential for better convergence and remedying potential perceptual loss imperfections. When the distances are reasonably large but satisfy the perceptual loss with high confidence, it will assign larger weights that counteract the color kernel. Only when these distances are extremely large do they become prohibitive, which is ideal. Hence, our optimizable filter is:
$\forall\mathcal{I}:\;OF_{\mathcal{K}^{*},\mathcal{R},\sigma_{c}}(\mathcal{I})=\underset{m,n}{\large\mbox{$\sum$}}PE_{\tiny\mbox{$M{\times}N$}}(\mathcal{I})\odot\mathcal{V}^{\mathcal{R},\mathcal{K}^{*},\sigma_{c}}$
$\mathcal{R}$ is a reference image for obtaining the color kernels, always chosen as $\mathcal{R}=x_{w}$. Next, we show how to learn $\mathcal{K}^{*}$’s weights.
5.2.3. Adversarial Filtering

### v2

@@ 5.1 Attacking Non-Semantic Watermarks
Final High-Frequency Destruction Procedure. (for non-semantic watermarks) Using the above components, we implement eq. (1). We only need to identify the rules for updating the sample to define our search space. Since we seek high-frequency disruptions that affect the rates at which individual pixels vary w.r.t. their neighbors, we should modify each pixel directly at each step. Thus, our algorithm accepts watermarked image $x_{w}$, LPIPS loss $\ell_{p}$, LPIPS threshold $t_{\ell_{p}}$, and $\|\|_{2}$ threshold $t_{\|\|_{2}}$, and finds modifier $\delta$ that maximizes DFL as follows:
[…]
$ReLU_{\ell}(x,\;y)$ denotes $max\{\ell(x,\;y)-t_{\ell},0\}$ for any loss $\ell$. Effectively, we instantiate $\ell_{d}$ from eq. (1) with two components: LPIPS loss $\ell_{p}$ and norm $\|\|_{2}$, each with its threshold, while DFL is the spectral loss. The vector $\delta$ captures the changes to $x_{w}$ to generate the non-watermarked $x_{nw}\!\!\!=\!\!x_{w}\!\!+\!\!\delta$. We minimize -DFL, maximizing DFL. The constants $c_{\ell_{p}}$ and $c_{\|\|_{2}}$ balance the spectral and perceptual losses. We enforce that each perceptual loss ($\ell_{d}$) component ($\ell_{p}$ and $\|\|_{2}$) remains below its threshold ($t_{\ell_{p}}\!$ or $t_{\|\|_{2}}$) through the term $ReLU_{\ell_{d}}(x_{w}\!\!+\delta,\;x_{w})$: When $\ell_{d}$ exceeds $t_{\ell_{d}}$, it evaluates to $\ell_{d}(x_{w}\!+\delta,\;x_{w})-t_{\ell_{d}}$, causing $\ell_{d}$’s gradients to produce a sample closer to $x_{w}$ in the next iteration. Otherwise, it returns 00, discarding $\ell_{d}$’s gradients and maximizing DFL unconditionally. $c_{\|\|_{2}}$ is a fixed large coefficient that strongly prevents the abovementioned peaks. $c_{\ell_{p}}$ is determined via a binary search: We start with a large $c_{\ell_{p}}$, assigning a larger weight to $\ell_{p}$’s gradients and run the optimization for several steps. If a solution is found, $c_{\ell_{p}}$ is decreased, increasing DFL’s dominance at the expense of larger visual deviations. Otherwise, $c_{\ell_{p}}$ is amplified. See Appendix A for details.
5.2 Attacking Semantic Watermarks
Semantic watermarks are in low frequencies, making them difficult to address. In addition to their large magnitudes (see §4.1.2), low frequencies correspond to many neighboring pixels varying gradually, indicating these watermarks are rooted in the image’s texture. Thus, removing them requires systemic and structural changes. Direct optimizations similar to §5.1 fail since they operate on the image and, at each step, individual pixels are updated in different directions. This targets higher frequencies, introducing significant changes yet only to a few pixels, failing to progress due to the perceptual constraints and incurring minimal changes at low frequencies. We could limit the procedure to optimize DFL over low-frequencies only. Yet, as significant changes are required in this range, without a methodical scheme, we will converge to a sub-optimal solution wherein a certain region receives substantial modifications, leading the perceptual loss to prevent further updates. This is because the process fails to account for inter-pixel links, and DFL does not spread the changes over the image.
(a) Basic: To calculate the output at $(i,\;j)$, the patch $PE^{i,j}_{\text{\tiny\mbox{$M\!{\times}\!N$}}}$ (where $M\!{\times}\!N$ is the kernel size) around $(i,\;j)$ is element-wise multiplied by the spectral kernel $\mathcal{K}$, summing over this product.
(b) Guided: The color kernel $\mathcal{C}^{\mathcal{I},\sigma\!_{{}_{c}}}_{i,j}$ is first computed, assigning a modulation factor to each neighbor in the vicinity $PE^{i,j}_{\text{{\tiny\mbox{$M\!{\times}\!N$}}}}$ of $(i,\;j)$. These are element-wise multiplied by the spectral kernel $\mathcal{K}$ to limit the contribution from neighbors whose values differ significantly from $(i,\;j)$’s. This yields the effective kernel $\mathcal{V}^{\mathcal{I},\mathcal{K},\sigma\!_{{}_{c}}}_{i,j}$ used to calculate the output at $(i,\;j)$ similar to basic filters.
Figure 3: Operation of known filters.
We develop a visually-aware attack inspired by image filtering. While filters apply systemic spectral disruptions, they fail to remove watermarks due to their content-agnostic nature, limiting the distortions they may introduce without leaving crucial visual cues. Our adversarial alternative retains quality by learning the filters’ weights to solve an optimization problem that maximizes low-frequency disruptions, resorting to a perceptual loss to preserve critical information. First, we motivate our design through guiding principles. We then provide technical details and present our optimizable filters. Finally, we introduce our attack.
5.2.1 Motivation
Convolution in space translates to multiplication in the frequency domain: Given the spectra of the image and filter, it produces an image with a spectrum equivalent to their product. Thus, filters are good candidates for spectral manipulations. Yet, the outputs at any pixel are calculated using the same kernel. Hence, their potential to remove watermarks is limited [23, 35] due to their content-agnostic nature, making them only applicable with mild parameters as stronger blurring impacts quality.
Guided Filtering: These advanced filters tackle the information loss under stronger distortions, allowing more effective noise reduction while preserving crucial data. They are edge-preserving, minimizing the disruptions at these critical boundary points between different objects. Instead of using a uniform kernel, they involve an additional per-pixel color kernel that weakens the filter at edge points based on geometric principles: If some values in the pixel’s vicinity drastically differ from it, then it lies on an edge, as these neighbors belong to other components and the spectral kernel’s strength here must be diminished to preserve defining information. The color kernel computes a modulation factor for every neighbor in the filter’s vicinity around $(i,\;j)$ based on its value’s difference from $(i,\;j)$’s. The effective kernel for calculating $(i,\;j)$’s output is the Hadamard product of the two kernels— see Fig. 3(b). Guided filters fail to remove watermarks despite their stronger blurring (see §6.2.1) since they merely follow geometric heuristics that fail to fully capture human vision or target specific image characteristics, limiting the allowed distortion. In practice, less visually-critical areas can be changed drastically, regardless.
5.2.2 Formalizing Filters
Given image  Q×Rsuperscript $\mathcal{I}\!\!\in\!\!\mathbb{R}^{\text{\tiny{ \mbox{$Q\!{\times}\!R$}}}}$, $PE^{i,j}_{\text{\tiny\mbox{$M\!{\times}\!N$}}}:\mathbb{R}^{\text{\tiny{ \mbox{$Q\!{\times}\!R$}}}}$  M×Nabsentsuperscript $\!\!\longrightarrow\!\!\mathbb{R}^{\text{\tiny{ \mbox{$M\!{\times}\!N$}}}}$ denotes the local patch extractor that outputs the $M{\times}N$ vicinity around its $(i,\;j)^{th}$ pixel s.t. this pixel is at the center (i.e., $PE^{i,j}_{\text{\tiny\mbox{$M\!{\times}\!N$}}}(\mathcal{I})(\text{\small\mbox{$\left\lfloor\frac{M}{2}\right\rfloor$}},\;\text{\small\mbox{$\left\lfloor\frac{N}{2}\right\rfloor$}})=\mathcal{I}(i,\;j)$)— see Fig. 3(a). The global patch extractor $PE_{\text{\tiny\mbox{$M\!{\times}\!N$}}}:\mathbb{R}^{\text{\tiny{ \mbox{$Q{\times}R$}}}}\!\!\longrightarrow\!\!\mathbb{R}^{\text{\tiny{ \mbox{$Q\!{\times}\!R{\times}M\!{\times}\!N$}}}}$ outputs a matrix that stores in cell $(i,\;j)$ the local extractor $PE^{i,j}_{\text{\tiny\mbox{$M\!{\times}\!N$}}}$. That is, $PE_{\text{\tiny\mbox{$M\!{\times}\!N$}}}(\mathcal{I})$’s $(i,\;j)^{th}$ cell contains the $M{\times}N$ patch around $(i,\;j)$ in $\mathcal{I}$. We formalize filters to present our solution.
[…]
Guided Filtering: These employ additional per-pixel color kernels to modulate distortions at critical points (Fig. 3(b)). Given kernel size $M{\times}N$, the color kernel for each $(i,\;j)$ is the matrix whose $(m,\;n)$ cell is $\mathcal{C}^{\mathcal{I},\sigma\!_{{}_{c}}}_{i,j}(m,\;n)$ = $e^{\mbox{\small${-\frac{\|\mathcal{I}(i+m-\left\lfloor\frac{M}{2}\right\rfloor,\;j+n-\left\lfloor\frac{N}{2}\right\rfloor)-\mathcal{I}(i,\;j)\|_{1}}{2\sigma\!_{{}_{c}}^{2}}}$}}$. $\sigma\!_{{}_{c}}$ is a permissiveness controlling the modulation intensity. The $\|\|_{1}$ norm accounts for each pixel potentially having multiple color channels. For instance, RGB images associate each pixel with three values. The effective kernel for computing the output at $(i\;j)$ is then $\mathcal{V}^{\mathcal{I},\mathcal{K},\sigma\!_{{}_{c}}}_{i,j}=\text{\small\mbox{$\frac{\mathcal{C}^{\mathcal{I},\sigma\!_{{}_{c}}}_{i,j}\odot\mathcal{K}}{\underset{m,n}{\text{\large\mbox{$\sum$}}}\mathcal{C}^{\mathcal{I},\sigma\!_{{}_{c}}}_{i,j}\odot\mathcal{K}}$}}$. At the center, $\mathcal{C}^{\mathcal{I},\sigma\!_{{}_{c}}}_{i,j}(\left\lfloor\frac{M}{2}\right\rfloor,\;\left\lfloor\frac{N}{2}\right\rfloor)=1$, indicating the contribution of the original $\mathcal{I}(i,\;j)$ is not decayed. For neighbors, large distances decrease their coefficients and limit their contributions. If we define $\mathcal{V}^{I,\mathcal{K},\sigma\!_{{}_{c}}}$ as the matrix whose $(i,\;j)^{th}$ cell contains $\mathcal{V}^{\mathcal{I},\mathcal{K},\sigma\!_{{}_{c}}}_{i,j}$, the guided filter is given as $GF_{\mathcal{K},\sigma_{c}}(\mathcal{I})=\underset{m,n}{\text{\large\mbox{$\sum$}}}PE_{\text{\tiny\mbox{$M\!{\times}\!N$}}}(\mathcal{I})\odot\mathcal{V}^{\mathcal{I},\mathcal{K},\sigma\!_{{}_{c}}}$.
Optimizable Filters (Ours): Guided filters’ drawback is that they employ the same spectral kernel $\mathcal{K}$ for all pixels, changing only the color kernel computed via the same formula at all locations. Instead, we allow each pixel $(i,\;j)$ to have a unique spectral kernel ${\mathcal{K}\!\!\!_{\;\;i,j}}$, with weights that can differ significantly between locations based on feedback from a perceptual loss that determines their visual importance and allows maximizing the destruction at less critical regions. Our filters consist of a set $\mathcal{K}^{*}$ of local ${\mathcal{K}\!\!\!_{\;\;i,j}}$’s. We include color kernels to avoid exploring unwanted directions that violate vital geometric constraints. While these may be limiting, ignoring them causes unnatural distortions if the differences between neighboring pixels are extremely large, making them essential for remedying potential perceptual loss imperfections. When the distances are reasonably large but satisfy the perceptual loss with high confidence, it will assign larger weights that counteract the color kernel. Only when these distances are extremely large do they become prohibitive, which is ideal. Our optimizable filter is:
$\forall\mathcal{I}:\;OF_{\mathcal{K}^{*},\mathcal{R},\sigma\!_{{}_{c}}}(\mathcal{I})=\underset{m,n}{\text{\large\mbox{$\sum$}}}PE_{\text{\tiny\mbox{$M{\times}N$}}}(\mathcal{I})\odot\mathcal{V}^{\mathcal{R},\mathcal{K}^{*},\sigma\!_{{}_{c}}}$
$\mathcal{R}$ is a reference image for obtaining the color kernels, set to $\mathcal{R}\equiv x_{w}$. We now show how to learn $\mathcal{K}^{*}$’s weights.
5.2.3 Adversarial Filtering
@@
where $\theta_{\mathcal{K}^{*}\!\!\!_{{}_{t}}}$ are the kernels’ weights and $\delta$ is a modifier that directly optimizes $x_{w}$, similar to eq. (2). We maximize the spectral loss by sequentially propagating the input through a network of filters. Recall that direct modifications are not effective for disrupting low frequencies in the absence of filters due to their inclination to cause high-frequency disturbances rather than systemic shifts in the values of many neighboring pixels in a gradual manner— see §5.2. Thanks to their convolutional nature, this no longer holds with our filters, making their outputs combinations of neighboring pixels. We still include $\delta$ that directly manipulates the input despite its inefficacy in the absence of filters. This is because the originally high-frequency disturbances caused by such modifications will now be distributed over neighbors of the receiving pixels due to our filters, translating them into useful low-frequency noise, causing further disruptions beyond the filters’ independent ability. We employ multiple ($T$) filters stacked after one another. The reason is that chaining several filters with various kernel sizes (geometries) targets different directions and slopes where the watermark could have been injected since watermarking is content-dependent and not all images can be changed similarly.
5.2.4 Spectral Loss
The former DFL is unsuitable for low frequencies (see §5.2) as it fails to force the changes to spread throughout the image. The collective energy at higher frequencies is much smaller in comparison, significantly reducing the total number of pixels that need to change to erase watermarks in these bands, which is achievable without violating the perceptual constraints. In contrast, sufficiently manipulating lower frequencies can rapidly change the structure, hindering convergence to an optimal solution if the distortion is not carefully distributed. This loss should also ensure our kernels perform the filtering operation, which DFL fails to enforce. Better alternatives are:
Filter Regularization Loss (FRL): It enforces that our kernels indeed operate as (smoothing) filters and is given as:
$&\forall x\in\mathbb{R}^{\text{\tiny{ \mbox{$Q\!{\times}\!R$}}}}:\;\textit{\text{FRL}}(x)=-\sum_{t\in[\![T]\!]}\textit{\text{FRL}}_{{}_{{\boldsymbol{t}}}}(x)\\[-5.0pt] &\;\;\;\;\;\;\;\;\;\;\;s.t.\;\forall t:\;\;\textit{\text{FRL}}_{{}_{{\boldsymbol{t}}}}(x)=\|x-{{{{{\mu}}}}\!_{{}_{{}_{\frac{1}{2}}}}\!\!\!\!\!\!\!}^{{}^{{}^{\text{\tiny\mbox{$M_{t}\!{\times}\!N_{t}$}}}}}\!\!\!\!\!\!(x)\|_{2}\end{aligned}$
$M_{t}{\times}N_{t}$ is the size of our $t^{th}$ kernel ${\mathcal{K}^{*}\!\!\!_{{}_{t}}}$ and ${{{{{\mu}}}}\!_{{}_{{}_{\frac{1}{2}}}}\!\!\!\!\!\!\!}^{{}^{{}^{\text{\tiny\mbox{$M_{t}\!{\times}\!N_{t}$}}}}}\!\!\!\!\!\!$ computes the median of this vicinity around each pixel. Maximizing FRL minimizes $\textit{\text{FRL}}_{{}_{{\boldsymbol{t}}}}$ for every filter $t$, which is the sum of all pixels’ distances from the median in the filter’s environment around each pixel, bringing pixels closer to these local medians. This eliminates variations to the extent possible under the perceptual loss, leading to smoother outputs. As the watermark is distributed over the image, such patches will exhibit gradual “watermarking” variations that will now be lost. A major benefit of FRL is output consistency.
Mean Pool Loss (MPL): FRL is insufficient, ensuring only proximity of the pixels to the local medians but imposing no constraints on these medians. The means of the various patches in the watermarked image are highly influenced by the watermark that causes many pixels to change their values. Due to the perceptual loss that strives to retain information, the procedure may converge to a sub-optimal solution, skewing these medians toward the original means and only slightly affecting pixel intensities, retaining many shifts associated with the watermark. Thus, we require MPL:
$\forall x,y\in\mathbb{R}^{\text{\tiny{ \mbox{$Q\!{\times}\!R$}}}}:\;\;\textit{\text{MPL}}^{{}^{\text{\tiny\mbox{$L\!{\times}\!P$}}}}(x,\;y)=\|\overline{x}^{{}^{\text{\tiny\mbox{$L\!{\times}\!P$}}}}\!\!\!-\overline{y}^{{}^{\text{\tiny\mbox{$L\!{\times}\!P$}}}}\|_{1}$
$\overline{x}^{{}^{\text{\tiny\mbox{$L\!{\times}\!P$}}}}\!\!$ denotes mean pooling with an $L{\times}P$ window. MPL takes the means over each $L{\times}P$ patch in both inputs and calculates the $\ell_{1}$ difference. As most pixels do not encode watermark information, maximizing the means’ differences (w.r.t. the watermarked image) while demanding proximity to the medians (using FRL) primarily affects pixels located far (those that encode the watermark) from all others, bringing them closer to the rest and eliminating watermarking variations. A proper low-frequency loss must spread the changes over multiple regions (see §5.2.4). MPL does so by dividing the image into patches and demanding that each receive a portion of the change, implying the number of patches must be large. Yet, each must contain enough pixels to manipulate low frequencies. We choose ${\text{\small\mbox{{MPL}}}}^{{}^{\text{\tiny\mbox{$5\!{\times}\!5$}}}}\!\!$ as a compromise, balancing these requirements. Calculating MPL for all filter sizes similar to FRL is less desirable as when the sizes shift away from this sweet spot, one of the conditions is violated. Note that, without FRL, MPL is insufficient as it can lead to all pixels in each patch receiving a similar modification $\epsilon$, simply changing its brightness. This satisfies the perceptual constraints but leaves the spectrum unaffected as the pixel variation rates remain identical.
5.2.5 Final Low-Frequency Destruction Procedure